# Enhanced Detection of Tiny Objects in Aerial Images

### v1

@@
Abstract–While one-stage detectors like YOLOv8 offer fast training speed, they often under-perform on detecting small objects as a trade-off. This becomes even more critical when detecting tiny objects in aerial imagery due to low-resolution targets and cluttered backgrounds. To address this, we introduce three enhancement strategies–input image resolution adjustment, data-augmentation and attention mechanism–that can be easily implemented on YOLOv8. We demonstrate that image size enlargement and the proper use of augmentation can lead to enhancement. Additionally, we designed a Mixture of Orthogonal Neural-modules Network(MoonNet) pipeline which consists of attention-augmented CNNs. Two well-known attention modules Squeeze-and-Excitation Block (SE Block) and Convolutional Block Attention Modules (CBAM) were integrated to the backbone of YOLOv8 with the increased number of channels and the MoonNet backbone obtained improved detection accuracy compared to that of the original YOLOv8. MoonNet further proved its adaptability and potential by achieving state-of-the-art performance on tiny object benchmark when integrated with the YOLC model. Our codes are available at:https://github.com/Kihyun11/MoonNet
@@ 1 Introduction
Advances in deep learning have transformed object detection, evolving from early sliding-window CNNs such as R-CNN [1] to modern pyramidal, end-to-end architectures. Two-stage detectors—including Faster R-CNN [2], Feature Pyramid Networks (FPN) [3], and Mask R-CNN [4]—first generate region proposals and then refine classification and localization, achieving state-of-the-art accuracy. In parallel, one-stage models prioritise speed by eliminating the proposal stage; seminal examples include SSD [5] and the “You Only Look Once” (YOLO) family [6] [7]. Although one-stage detectors excel in training speed, they tend to struggle with detecting tiny objects as a design trade-off. This limitation is especially pronounced in aerial imagery, where targets often occupy only a few pixels amid cluttered backgrounds. This work addresses the challenge of tiny-object detection, using YOLOv8. Our contributions are threefold: (i) higher input resolution, (ii) targeted data augmentation, and (iii) an attention-augmented backbone—each yielding measurable gains.
@@ 2 Related Work
Previous studies using YOLO: Plenty of researches specifically targeting the enhancement of tiny object detection using the YOLO model already exists. Ma et al. [8] improve YOLOv8s for tiny-object detection by substituting standard strided convolutions with the SPD-Conv module and replacing the PANet neck with SPANet, achieving finer feature preservation and stronger multi-scale fusion. Hu et al. [9] enhance YOLOv5 through an Efficient Spatial Pyramid Pooling (ESPP) module, raising small-object accuracy in aerial imagery while keeping the network lightweight.
Tiny object detection in general: Prior work has explored various methods for tiny object detection. Xu et al. [10] suggests that effective enhancement methods for tiny object detection can be categorized into five: multi-scale feature learning, context based detection, data augmentation, designing better training strategy and label assignment strategy. High resolution imagery is crucial as well and numerous approaches were made to improve the resolution of the images such as QueryDet [11]. Some studies such as YOLC [12] even use cropping-method to make the training more efficient.
@@
Attention Mechanism: Hu et al. [15] proposed the Squeeze-and-Excitation (SE) block to enhance CNNs by modeling inter-channel dependencies and recalibrating feature responses. Woo et al. [16] extended this with the Convolutional Block Attention Module (CBAM), which combines channel and spatial attention to better focus on informative regions—an advantage in tiny object detection where spatial precision is crucial.
@@ 3.1 Enhancement Strategy
In total, three different enhancement strategies were implemented on the YOLOv8 to obtain efficient training.
Varying the input image resolution: Three different input image resolutions 640, 800 and 928 were tested.
Data Augmentation: A diverse set of data augmentation techniques was applied. These included geometric transformations (rotations, horizontal and vertical flips), photometric adjustments (brightness and contrast variation), and random noise injection. While these augmentations simulated real-world variations, the primary motivation was to mitigate class imbalance—particularly for under-represented tiny object categories.
Attention Modules Integration: To improve the model’s sensitivity to fine-grained details, attention mechanisms were integrated into the YOLOv8 backbone. Specifically, Squeeze-and-Excitation (SE) blocks and Convolutional Block Attention Modules (CBAM) were inserted directly into select convolutional layers of the backbone. Instead of implementing them as standalone modules, the convolution layer class was extended to perform channel-wise recalibration (SE) and spatial attention (CBAM) inline. This design minimized architectural complexity and computational overhead while enhancing the model’s focus on salient regions critical for tiny object detection.
@@ 3.2 Final Configuration
Following a multi-stage optimization process, the final model configuration was determined by selecting the most effective combinations of image resolution, data augmentation techniques, and network architecture. Using the modified backbone with integrated attention modules and the best-performing input image resolution setting, the final model was trained from scratch.
3.3 Integration with YOLC
The attention modules are also integrated to the backbone of the YOLC[12] with the following modifications:
Squeeze-Excitation Operation (Ours)
$x$ $\in\mathbb{R}^{C\times H\times W},$ (1a)
$c$ $=\mathrm{GlobalAvgPool}(x)\in\mathbb{R}^{C\times 1\times 1},$ (1b)
$z$ $=W_{2}\phi(W_{1}c),\qquad\phi=\mathrm{ReLU},$ (1c)
$y_{\text{original}}$ $=x\odot\sigma(z),\qquad\sigma=\mathrm{sigmoid},$ (1d)
$y_{\text{ours}}$ $=x\odot\bigl(1+\tanh(z)\bigr).$ (1e)
CBAM Operation (Ours)
$x$ $\in\mathbb{R}^{C\times H\times W},$ (2a)
$c_{\mathrm{avg}}$ $=\mathrm{GlobalAvgPool}(x),$ (2b)
$c_{\mathrm{max}}$ $=\mathrm{GlobalMaxPool}(x),$ (2c)
$z_{c}$ $=W_{2}\,\phi(W_{1}c_{\mathrm{avg}})+W_{2}\,\phi(W_{1}c_{\mathrm{max}}),$ (2d)
$x_{c}^{\mathrm{original}}$ $=x\odot\sigma(z_{c}),$ (2e)
$x_{c}^{\mathrm{ours}}$ $=x\odot\bigl(1+\tanh(z_{c})\bigr),$ (2f)
$f_{\mathrm{avg}}$ $=\mathrm{Avg}_{\text{chan}}(x_{c}^{\bullet}),\quad f_{\mathrm{max}}=\mathrm{Max}_{\text{chan}}(x_{c}^{\bullet}),$ (2g)
$z_{s}$ $=\mathrm{Conv}_{k\times k}\!\big([\,f_{\mathrm{avg}},\,f_{\mathrm{max}}\,]\big),$ (2h)
$y_{\mathrm{original}}$ $=x_{c}^{\mathrm{original}}\odot\sigma(z_{s}),$ (2i)
$y_{\mathrm{ours}}$ $=x_{c}^{\mathrm{ours}}\odot\bigl(1+\tanh(z_{s})\bigr).$ (2j)
Parameters & notes:  $W_{1}\!\in\!\mathbb{R}^{m\times C},\;W_{2}\!\in\!\mathbb{R}^{C\times m},\;\\ m=\max\!\bigl(8,\lfloor C/r\rfloor\bigr)\ \text{for reduction ratio }r>0$. Channel logits $z$ (SE) and $z_{c}$ (CBAM) lie in $\mathbb{R}^{C}$ and are broadcast over $H\times W$ (equivalently reshape to $C\times 1\times 1$). Spatial logits $z_{s}\in\mathbb{R}^{1\times H\times W}$ are broadcast over channels. Here $\odot$ denotes element-wise multiplication, and $x_{c}^{\bullet}$ denotes the branch (original/ours) passed to the spatial gate. $x$ denotes the input feature map and the subscript chanchan means channel-wise. All the variables with ‘original’ subscript denotes the original operation whereas that with ‘ours’ denotes the modified operation.
Motivation. In preliminary trials, with the attention-attached HRNet backbone, the YOLC model learned nothing. We claimed the major cause of this result as a sigmoid gating which halves the output signal right away. Our modification replaces the sigmoid gate with a residual, zero-centered gate $x\odot\bigl(1+\tanh(\cdot)\bigr)$ and uses identity-safe init (set the last projection(s) to zero) so $z{=}0\Rightarrow y_{\text{ours}}{=}x$ at step 0.
Placement.  Attention is applied per branch at the outputs of the final HRNet stage (post-branch adapters), not inside internal convolutional blocks. Let $\{f^{(i)}\}_{i=1}^{B}$ be the multi-resolution maps (B=4B{=}4); we compute $y^{(i)}=\mathcal{A}^{(i)}(f^{(i)})$, $\mathcal{A}^{(i)}\!\in\!\{\text{SE BLOCK},\text{CBAM}\}$. Because gating is multiplicative and zero-initialized, tensor shapes and channels are unchanged, so the YOLC neck/head need no code changes.
@@ 4.1 Datasets
Modified DOTAv2.0. Given the focus of this research on enhancing the detection of tiny objects, a dataset tailored to such targets was essential. To this end, we modified the original DOTA 2.0 dataset by selectively retaining only those object categories that consistently exhibit small spatial footprints in aerial imagery. The resulting modified dataset comprises five classes: small vehicle, large vehicle, ship, plane, and storage tank. Since the modified dataset was used, the evaluation could not be done through evaluation server but was performed locally using the test set. The dataset is in the repository.
Table 1: Object categories of the Modified-DOTA
@@
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 2: Dataset Size Comparison
Dataset Average pixel size
Aerial Imagery Dataset
UCAS-AOD [17] $\approx 4149\,\text{px}^{2}\ (64.4\times 64.4)$
VisDrone [18] $\approx 2490\,\text{px}^{2}\ (49.9\times 49.9)$
DOTAv2.0 [19] $\approx 3394\,\text{px}^{2}\ (58.2\times 58.2)$
Modified-DOTA (Ours) $\approx\textbf{968}\,\textbf{px}^{2}\ \textbf{(31.1}\times\textbf{31.1)}$
@@
Modified-DOTA (Ours) $\approx\textbf{968}\,\textbf{px}^{2}\ \textbf{(31.1}\times\textbf{31.1)}$
As it can be seen in the table 2, the retained objects in our modified dataset occupies approximately 968 pixels² corresponding roughly to a 31×31 pixel bounding region.This area is below widely accepted thresholds for small objects, such as the 1024 pixels² cutoff defined in the COCO dataset for “small” instances [20].
VisDrone. We also evaluated on the VisDrone benchmark to assess generalization and enable fair comparison. Table 5 reports backbone ablations (identifying the best backbone), and Table 8 benchmarks our approach against state-of-the-art methods in tiny-object detection. Since the VisDrone server is currently unavailable, we used the validation set for evaluation.
@@ 4.2 Training Setting
All training was conducted in a secure cloud using one NVIDIA L40S GPU, 32 vCPUs, and 125 GB RAM. The environment ran PyTorch 2.2.0, Python 3.10, and CUDA 12.1.1 on Ubuntu 22.04.
Trial Screening: Input image adjustment, augmentation, and backbone testing were performed over 50 epochs (batch size 4, 640×640 image size) using the SGD optimizer and a modified, non-augmented DOTA dataset. The trainings used pretrained YOLOv8n.
Final Training: The YOLOv8n with the default and MoonNet backbones were trained for 150 epochs with batch size 4 and the selected candidates setting. The YOLC model trainings were conducted for 50 epochs.
@@ 4.3 Input image resolution adjustment
Identified 928 as yielding the highest performance.
@@
($\times640$) 0.323 0.180 0.279 0.638
($\times800$) 0.344 0.199 0.301 0.741
($\times928$) ✓ 0.372 0.208 0.331 0.635
@@ 4.4 Data Augmentation
A total three of different augmentation packages were tested and the implemented augmentation types were explained in table 4. Geo-denotes geometric transformations, BBox-denotes bounding box transformation, and Color/Blur-denotes noise and color adjustment.
@@
Package Geo BBox Color/Blur $_{50} APAP$ Recall Precision
Ver 1 No Augmentation 0.323 0.180 0.279 0.638
Ver 2 0.371 0.213 0.325 0.662
Ver 3 0.344 0.194 0.303 0.648
4.5 Attention Mechanism
Total six attention-augmented backbones were tested in table 5 and the backbone designs are displayed in table 6. The two modules were attached to the CNN layers of the original YOLO backbone in different arrangements.
MoonNet: The fifth design obtained the best performance among the backbone designs. We name this particular backbone with SE Block-CBAM arrangement as Mixture of Orthogonal Neural-modules Network(MoonNet) and this backbone is used for the final training adapted to the YOLOv8 and YOLC model.

### v2

@@
Abstract
While one-stage detectors like YOLOv8 offer fast training speed, they often under-perform on detecting small objects as a trade-off. This becomes even more critical when detecting tiny objects in aerial imagery due to low-resolution targets and cluttered backgrounds. To address this, we introduce four enhancement strategies–input image resolution adjustment, data-augmentation, attention mechanisms and an alternative gating function for attention modules–that can be easily implemented on YOLOv8. We demonstrate that image size enlargement and the proper use of augmentation can lead to enhancement. Additionally, we designed a Mixture of Orthogonal Neural-modules Network (MoonNet) pipeline which consists of multiple attention modules-augmented CNNs. Two well-known attention modules Squeeze-and-Excitation (SE) Block and Convolutional Block Attention Modules (CBAM) were integrated into the backbone of YOLOv8 to form MoonNet design and the MoonNet backbone obtained improved detection accuracy compared to the original YOLOv8 backbone and the single-type attention modules-augmented backbone. MoonNet further proved its adaptability and potential by achieving state-of-the-art performance on the tiny-object benchmark when integrated with the YOLC model. Our code is available at:https://github.com/Kihyun11/MoonNet
@@ 1 Introduction
Advances in deep learning have transformed object detection, evolving from early sliding-window CNNs such as R-CNN [1] to modern pyramidal, end-to-end architectures. Two-stage detectors—including Faster R-CNN [15], Feature Pyramid Networks (FPN) [8], and Mask R-CNN [2]—first generate region proposals and then refine classification and localization, achieving state-of-the-art accuracy. In parallel, one-stage models prioritize speed by eliminating the proposal stage; seminal examples include SSD [11] and the “You Only Look Once” (YOLO) family [13] [14]. Although one-stage detectors excel in training speed, they tend to struggle with detecting tiny objects as a design trade-off. This limitation is especially pronounced in aerial imagery, where targets often occupy only a few pixels amid cluttered backgrounds. Many studies have been conducted to tackle this problem and delivered significant enhancement using YOLO model family. Our research is highly influenced by both Zhu et al[22] and Kang et al[5] which successfully demonstrated the detection accuracy gain in aerial images using the attention modules. Kang et al. especially show that their multiple attention module augmented architecture is better than the SE Block-Only architecture. We further extended this idea to investigate whether the mixed use of attention modules can be always beneficial and what other factors can play an important role in detecting tiny objects using YOLOv8 and YOLC framework. Our contributions are fourfold: (i) effectiveness of higher input resolution, (ii) importance of targeted data augmentation, (iii) compare a hybrid attention-augmented backbone against the single-type attention augmented backbone and obtain the MoonNet design which out-performed among the design candidates, and (iv) suggest a possible alternative gating mechanism for attention modules.
@@ 2 Related Work
Tiny object detection in general: Prior work has explored various methods for tiny object detection. Xu et al. [18] suggest that effective enhancement methods for tiny object detection can be categorized into five: multi-scale feature learning, context-based detection, data augmentation, designing better training strategy and label assignment strategy. High resolution imagery is crucial as well and numerous approaches were made to improve the resolution of the images such as QueryDet [20]. Some studies such as YOLC [10] even use cropping method to make the training more efficient.
@@
Attention Module & Gating Function: Hu et al. [3] proposed the Squeeze-and-Excitation (SE) block to enhance CNNs by modeling inter-channel dependencies and recalibrating feature responses. Woo et al. [17] extended this with the Convolutional Block Attention Module (CBAM), which combines channel and spatial attention to better focus on informative regions—an advantage in tiny object detection where spatial precision is crucial. Both modules use sigmoid as a gating function to allow the model to focus on relevant features and avoid vanishing gradients.
Previous studies using YOLO: Plenty of research specifically targeting the enhancement of tiny object detection using the YOLO model already exists. Ma et al. [12] improve YOLOv8s for tiny-object detection by substituting standard strided convolutions with the SPD-Conv module and replacing the PANet neck with SPANet, achieving finer feature preservation and stronger multi-scale fusion. Hu et al. [4] enhance YOLOv5 through an Efficient Spatial Pyramid Pooling (ESPP) module, raising small-object accuracy in aerial imagery while keeping the network lightweight.
Relation to prior work: Implementation of attention modules on the YOLO architecture had been explored previously. Zhu et al. [22] introduces TPH-YOLO, which demonstrates the effectiveness of the CBAM with YOLOv5 architecture when detecting objects in aerial imagery. Kang et al[5]’s SE-CBAM-YOLOv7 used both SE Block and CBAM modules on the YOLOv7’s architecture to detect tiny aircraft in aerial images. Unlike Zhu et al. and Kang et al., we focused more on testing several backbone designs with different attention module orientations to check the validity of using multiple attention modules and obtaining the best backbone candidate for enhancement.
We have also made an algorithmic modification on the SE Block and CBAM. As mentioned earlier, sigmoid is a commonly used gating function for the attention modules. A few other modules uses sigmoid as well, for instance, ECA-Module [16]. In this work, we tested a different gating mechanism $\bigl(1+\tanh(\cdot)\bigr)$,which is an identity safe gating that can preserve the original feature and also amplifies the scarce relevant features from tiny objects.
@@ 3.1 Enhancement Strategy
In total, four different enhancement strategies were implemented on YOLOv8 to obtain efficient training.
Varying the input image resolution: Three different input image resolutions 640, 800, and 928 were tested.
Data Augmentation: A total of three different sets of data augmentation techniques were applied. The first set is the default augmentation setting provided by the Ultralytics. The second set is geometric transformations-focused package and the third set is color space adjustments-focused package. While these augmentations simulated real-world variations, the primary motivation was to mitigate class imbalance—particularly for under-represented tiny object classes.
Attention Modules Integration: To improve the model’s sensitivity to fine-grained details, attention modules were integrated into the YOLOv8 backbone. Specifically, Squeeze-and-Excitation (SE) blocks and Convolutional Block Attention Modules (CBAM) were inserted directly into select convolutional layers of the backbone. Instead of implementing them as standalone modules, the convolution layer class was extended to perform channel-wise recalibration (SE) and spatial attention (CBAM) inline.
Identity-safe gating mechanism: While attention modules are implemented on the backbone, the gating function is switched. Thus, one set of two backbones with the same attention module arrangement but different gating was trained and tested. The modifications are presented in Equations1 and 2 and the full mathematical operations of SE Block and CBAM are available via Hu et al.[3] and Woo et al.[17]:
Squeeze-Excitation Operation (Ours)
$y_{\text{original}}$ $=x\odot\sigma(z),\qquad\sigma=\mathrm{sigmoid},$ (1a)
$y_{\text{ours}}$ $=x\odot\bigl(1+\tanh(z)\bigr).$ (1b)
CBAM Operation (Ours)
$x_{c}^{\mathrm{original}}$ $=x\odot\sigma(z_{c}),$ (2a)
$x_{c}^{\mathrm{ours}}$ $=x\odot\bigl(1+\tanh(z_{c})\bigr),$ (2b)
$y_{\mathrm{original}}$ $=x_{c}^{\mathrm{original}}\odot\sigma(z_{s}),$ (2c)
$y_{\mathrm{ours}}$ $=x_{c}^{\mathrm{ours}}\odot\bigl(1+\tanh(z_{s})\bigr).$ (2d)
Parameters & notes:  Channel logits $z$ (SE) and $z_{c}$ (CBAM) lie in $\mathbb{R}^{C}$ and are broadcast over $H\times W$ (equivalently reshape to $C\times 1\times 1$). Spatial logits $z_{s}\in\mathbb{R}^{1\times H\times W}$ are broadcast over channels. Here $\odot$ denotes element-wise multiplication. $x$ denotes the input feature map and the subscript $c$ means channel-wise. All variables with the original subscript denote the original operation whereas that with ours denotes the modified operation.
Placement.  For the YOLO-model, the modules were applied to the convolutional layers in the backbone. For the YOLC-model, the attention is applied per branch at the outputs of the final HRNet stage (post-branch adapters), not inside internal convolutional blocks. Let $\{f^{(i)}\}_{i=1}^{B}$ be the multi-resolution maps (B=4B{=}4); we compute $y^{(i)}=\mathcal{A}^{(i)}(f^{(i)})$, $\mathcal{A}^{(i)}\!\in\!\{\text{SE BLOCK},\text{CBAM}\}$. Because gating is multiplicative, tensor shapes and channels are unchanged, so the YOLC neck/head need no code changes.
@@ 3.2 Final Configuration
Following a multi-stage enhancement optimization process, the final model configuration was determined by selecting the most effective combinations of image resolution, data augmentation techniques, network architecture, and gating function. Using the modified backbone with integrated attention modules and the best-performing input image resolution setting, the final model was trained from scratch.
@@ 4.1 Datasets
Modified DOTAv2.0. We modified the original DOTA 2.0 dataset by selectively retaining only those object categories that consistently exhibit small spatial footprints in aerial imagery. The resulting modified dataset comprises five classes: small vehicle, large vehicle, ship, plane, and storage tank. Since the modified dataset was used, the evaluation could not be done through evaluation server but was performed locally using the validation set. The dataset is fully reproducible and the method is provided in the repository. Refer to Table 1, the retained objects occupy approximately 2214 pixels² (47×47).
VisDrone. We also evaluated on the VisDrone benchmark in Table 4 and Table 7 to assess generalization and enable fair comparison. Since the VisDrone server is currently unavailable, we used the validation set for evaluation.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fig. 1: Object categories of the Modified-DOTA
Table 1: Pixel size comparison
Object Classes Average pixel size
Small Vehicle $\approx 1458\,\text{px}^{2}\ (38\times 38)$
Large Vehicle $\approx 12407\,\text{px}^{2}\ (111\times 111)$
Ship $\approx 1199\,\text{px}^{2}\ (35\times 35)$
Plane $\approx 580\,\text{px}^{2}\ (24\times 24)$
Storage Tank  $\approx\,2591\,\text{px}^{2}\ (51\times 51)$
@@
Modified-DOTA (Ours) $\approx\textbf{2214}\,\textbf{px}^{2}\ \textbf{(47}\times\textbf{47)}$
@@ 4.2 Training setting
YOLOv8: All the training runs were conducted on a local workstation equipped with an NVIDIA RTX 4080 SUPER GPU (16 GB VRAM). The environment ran Ultralytics 8.3.239, Python 3.9.10, and PyTorch 2.5.1 on CUDA 12.1
YOLC: All the trainings were conducted in a secure cloud using one NVIDIA A40 GPU. The environment ran PyTorch 2.1.0, Python 3.10, and CUDA 11.8.0 on Ubuntu 22.04.
Trial screening: Input image adjustment, augmentation, and backbone testing were performed over 50 epochs (batch size 4, $640\times 640$ image size) using the AdamW optimizer and a modified, default-augmented DOTA dataset. The trainings used a pretrained YOLOv8n-obb model.
Final training: The YOLOv8n-obb with the default and MoonNet backbones were trained for 150 epochs with batch size 4 and the selected candidates setting. The YOLC model trainings were conducted for 50 epochs.
@@ 4.3 Input image resolution adjustment
We identified the image size of (928x928) as yielding the highest performance in $_{50}, APAP$, and Recall.
@@
($\times640$) 0.621 0.436 0.539 0.796
($\times800$) 0.665 0.489 0.583 0.821
($\times928$) ✓ 0.696 0.525 0.613 0.819
@@ 4.4 Data augmentation
A total of three of different augmentation packages were tested and the implemented augmentation types were explained in Table 3. Default denotes default augmentation setting provided by the Ultralytics, Geo denotes geometric transformations, and Color denotes HSV color space adjustment. The exact augmentations are provided in the repository.
@@
Package Default Geo Color $_{50} APAP$ Recall Precision
Ver 1 0.621 0.436 0.539 0.796
Ver 2 0.604 0.416 0.530 0.763
Ver 3 0.616 0.428 0.536 0.786
4.5 Attention & gating mechanism
A total of six attention-augmented backbones were tested in Table 4 and the designs are displayed in Figure 2. Dotted lines are used to distinguish between backbones employing single-type versus multiple-type modules. In the trainings using modified DOTA, the third design outperformed other candidates. The top 3 candidates were then trained with VisDrone dataset to check performance and the second design (Only-CBAM) achieved the highest accuracy.
The third design: Although the third design failed to obtain the best performance when trained using VisDrone, we decided to use the third design for the final training configuration since it edges out other backbones when trained with DOTA. We name this particular backbone with SE Block-CBAM arrangement as Mixture of Orthogonal Neural-modules Network(MoonNet).
Unfortunately, identity-safe gating $(1+\tanh(\cdot)\bigr)$ failed to achieve enhancement when trained with YOLO models. Thus, we did not run additional trainings using tanh gating when using VisDrone. However, tanh gate achieved enhancement when adapted to YOLC framework in Table 7 and 8.
Refer to caption $(640\times 640\times 3)$ $(640\times 640\times 3)$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption $(320\times 320\times 64\times w)$ $(320\times 320\times 64\times w)$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption $(160\times 160\times 128\times w)$ $(160\times 160\times 128\times w)$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption $(80\times 80\times 256\times w)$ $(80\times 80\times 256\times w)$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption $(40\times 40\times 512\times w)$ $(40\times 40\times 512\times w)$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption $(20\times 20\times 1024\times w)$ $(20\times 20\times 1024\times w)$
Fig. 2: Backbone designs